# labelinglabel

### v2

@@ 1. Introduction
Let $G$ be a countable discrete group, $X$ be a countable discrete set, $\alpha:G\curvearrowright X$ be an action, $A$ be a finite set, $\varphi:G\rightarrow\textrm{Sym}(A)$ be a map. For a finite subset $F\subseteq G$ and $\epsilon>0$, $\varphi$ is called $(F,\epsilon)$-multiplicative if $d(\varphi(gh),\varphi(g)\varphi(h))<\epsilon$ for all $g,h\in F$, where $d$ denotes the normalized Hamming distance. For finite subsets $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$, $\varphi$ is called an $(F,E,\epsilon)$-orbit approximation of $\alpha$ if there exists a finite set $B$ and a subset $S\subseteq A$ s.t. $|S|>(1-\epsilon)|A|$ and for each $s\in S$ there is an injective map $\pi_{s}:E\hookrightarrow B$ s.t. $\pi_{\varphi(g)s}(x)=\pi_{s}(\alpha(g^{-1})x)$ for all $s\in S$, $g\in F$, $x\in E$, whenever $\varphi(g)s\in S$ and $\alpha(g^{-1})x\in E$.
@@
$\alpha$ is called sofic if for all finite subsets $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$, there exists a finite set $A$ and a map $\varphi:G\rightarrow\textrm{Sym}(A)$ which is unital, $(F,\epsilon)$-multiplicative, and an $(F,E,\epsilon)$-orbit approximation of $\alpha$.
@@
The first example is the left action of a sofic group $G$ on itself. We show (in Theorem 2.14 and Theorem 2.12) that this is sofic if and only if $G$ is sofic. More generally, we are able to show (in Theorem 2.14) that $G$ acting on the coset space $G/H$ is sofic where $H$ is any locally finite subgroup. On the other side, we have (see Theorem 2.12) that if the action of $G$ on a coset space $G/H$ is sofic, then there is a normal subgroup $N$ of $G$ such that $N\leq H$ and $G/N$ is sofic. This by itself could give help to finding new examples of sofic groups.  [^1] We would like to remind the readers that it remains a puzzling question whether $G/N$ is sofic for $G$ sofic and $N$ is a finite normal subgroup, say even $|N|=2$.
@@
The proof of the main result is very much inspired by and pushes the ideas of [HS18]. The above result also applies to the setting of amalgamated free generalized wreath product (see Theorem 3.7). The unfamiliar reader is directed to Definition 3.3. At this point we would like to mention that the natural variant of the above result can also be proved in the context of hyperlinearity, see Theorems 3.8, 3.9 and Corollary 3.10.
@@ 2. Sofic actions
Let $G$ be a countable discrete group, $X$ be a countable discrete set, $\alpha:G\curvearrowright X$ be an action, $A$ be a finite set, $\varphi:G\rightarrow\textrm{Sym}(A)$ be a map (not necessarily a homomorphism):
@@
    Recall that $G$ is called sofic if for all finite subsets $F\subseteq G$ and $\epsilon>0$ there exists a finite set $A$ and a map $\varphi:G\to\textrm{Sym}(A)$ which is unital, $(F,\epsilon)$-multiplicative, and $d(1,\varphi(g))>1-\epsilon$ for all $g\in F\setminus\{e\}.$
@@
    $\alpha$ is called sofic if for all finite subsets $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$, there exists a finite set $A$ and a map $\varphi:G\rightarrow\textrm{Sym}(A)$ which is unital, $(F,\epsilon)$-multiplicative, and an $(F,E,\epsilon)$-orbit approximation of $\alpha$.
@@
Let $[n]=\{1,\cdots,n\}$, $\mu_{n}$ be the normalized counting measure on $[n]$, $\mathcal{U}$ be a free ultrafilter on $\mathbb{N}$. Then the Loeb measure space, denoted by $\prod_{\mathcal{U}}([n],\mu_{n})$, is defined to have the underlying set $\prod_{\mathcal{U}}[n]$. For a sequence $(A_{n}\subseteq[n])$, we define $\mu_{\mathcal{U}}((A_{n})_{\mathcal{U}})=\lim_{\mathcal{U}}\mu_{n}(A_{n})$. This can be extended to a measure on the $\sigma$-algebra generated by all such $(A_{n})_{\mathcal{U}}$.
@@
Let $\mathcal{U}$ be a free ultrafilter on $\mathbb{N}$. For each $n$, let $X_{n}$ be the collection of all maps $[n]\rightarrow\mathbb{N}$ and $d_{n}$ be the normalized Hamming distance on $X_{n}$. Let $\mathbb{X}^{2}_{\mathcal{U}}=\prod_{\mathcal{U}}(X_{n},d_{n})$. We define the following action $\mathbb{S}^{2}_{\mathcal{U}}$ of the universal sofic group $\prod_{\mathcal{U}}(S_{n},d)$ on $\mathbb{X}^{2}_{\mathcal{U}}$ by $\mathbb{S}^{2}_{\mathcal{U}}((g_{n})_{\mathcal{U}})((f_{n})_{\mathcal{U}})=(f_{n}\circ g_{n}^{-1})_{\mathcal{U}}$. We easily observe that this is a well-defined action that preserves the metric $d_{\mathcal{U}}$. We shall call $\mathbb{S}^{2}_{\mathcal{U}}$ the second universal sofic action.
@@
()($\Rightarrow$)( ⇒ ) Fix increasing sequences of finite subsets $F_{1}\subseteq F_{2}\subseteq\cdots\subseteq G$ and $E_{1}\subseteq E_{2}\subseteq\cdots\subseteq X$ s.t. $\cup_{i}F_{i}=G$, $\cup_{i}E_{i}=X$. Fix a decreasing sequence $\epsilon_{i}>0$ s.t. $\lim_{i}\epsilon_{i}=0$. For each $i$, let $\varphi_{i}:G\rightarrow\textrm{Sym}(A_{i})$ be a unital, $(F_{i},\epsilon_{i})$-multiplicative, and an $(F_{i},E_{i},\epsilon_{i})$-orbit approximation of $\alpha$. By taking the Cartesian products of $A_{i}$ with auxiliary finite sets if necessary, we may assume $|A_{i}|$ is strictly increasing. By definition of $(F_{i},E_{i},\epsilon_{i})$-orbit approximation, there exists a finite set $B_{i}$ and a subset $S_{i}\subseteq A_{i}$ s.t. $|S_{i}|>(1-\epsilon_{i})|A_{i}|$ and for each $s\in S_{i}$ there is an injective map $\pi^{i}_{s}:E_{i}\hookrightarrow B_{i}$ s.t. $\pi^{i}_{\varphi_{i}(g)s}(x)=\pi^{i}_{s}(\alpha(g^{-1})x)$ for all $s\in S_{i}$, $g\in F_{i}$, $x\in E_{i}$, whenever $\varphi_{i}(g)s\in S_{i}$ and $\alpha(g^{-1})x\in E_{i}$. By embedding $B_{i}$ into $\mathbb{N}$ we may take the co-domain of $\pi^{i}_{s}$ to be $\mathbb{N}$. Let $\mathcal{U}$ be any free ultrafilter on $\mathbb{N}$ containing the set $\{|A_{i}|:i\geq 1\}$. We may then define $\varphi:G\rightarrow\prod_{\mathcal{U}}(S_{n},d)$ by defining $\varphi(g)=(g_{n})_{\mathcal{U}}$ with $g_{n}=\varphi_{i}(g)$ whenever $n=|A_{i}|$ and $g_{n}=1$ otherwise. Since $\{|A_{i}|:i\geq 1\}\in\mathcal{U}$ and $\varphi_{i}$ is $(F_{i},\epsilon_{i})$-multiplicative, we see that $\varphi$ is a group homomorphism.
@@
()($\Leftarrow$)( ⇐ ) For each $g\in G$, we shall write $\varphi(g)=(g_{n})_{\mathcal{U}}$. Since $\varphi(1_{G})=1$, we shall in particular choose $\varphi(1_{G})=(1)_{\mathcal{U}}$. Let $\varphi_{n}:G\rightarrow S_{n}$ be defined by $\varphi_{n}(g)=g_{n}$. Now, fix finite $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$, we shall show that there exists $n$ s.t. $\varphi_{n}$ is unital, $(F,\epsilon)$-multiplicative, and an $(F,E,\epsilon)$-orbit approximation of $\alpha$. We observe that $\varphi_{n}$ is unital for all $n$. Since $F$ is finite and $\varphi$ is a group homomorphism, there exists $L_{1}\in\mathcal{U}$ s.t. for all $n\in L_{1}$, $\varphi_{n}$ is $(F,\epsilon)$-multiplicative.
@@
[…]
@@
Since $G$ is sofic, there exists a finite set $A$ and a map $\varphi:G\rightarrow\textrm{Sym}(A)$ which is unital and $(F^{\prime},\epsilon^{\prime})$-multiplicative and satisfies $d(1,\varphi(g))>1-\epsilon^{\prime}$ for all non-identity $g\in F^{\prime}$.
Fix $g,h\in F$, $g\neq h$, then, since $F$ is symmetric and since $\varphi$ is unital and $(F^{\prime},\epsilon^{\prime})$-multiplicative, we have,
@@
Hence, |$|\{s\in A:\varphi(g)s\neq\varphi(h)s\}|=|\{s\in A:s\neq\varphi(g)^{-1}\varphi(h)s\}|>(1-3\epsilon^{\prime})|A|$. So,
@@
We also observe that, for any fixed $g,h\in F$, |$|\{s\in A:\varphi(gh)s=\varphi(g)\varphi(h)s\}|>(1-\epsilon^{\prime})|A|$. Therefore,
@@
Let $F\subseteq G$, $E\subseteq G/N$ be finite subsets. Fix $\epsilon>0$. Let $q:G\rightarrow G/N$ be the natural quotient map. Fix $\sigma:G/N\rightarrow G$ an arbitrary section of $q$. Let $U=\{\sigma(x)^{-1}g\sigma(g^{-1}x):x\in E,g\in F\}$. We observe that $U$ is finite and $U\subseteq N$. Let $N^{\prime}=\langle U\rangle$. Since $N$ is locally finite, $N^{\prime}$ is a finite group. Let,
@@
We observe that since $N^{\prime}$ is a subgroup, $1_{G}\in N^{\prime}$, so $\sigma(E)^{-1}\subseteq N^{\prime}\cdot\sigma(E)^{-1}\subseteq F^{\prime}$. Let $\epsilon^{\prime}=\frac{\epsilon}{|E|+1}$. Since $F^{\prime}$ is finite, by Lemma 2.13, there is a finite set $A$ and a unital, $(F^{\prime},\epsilon^{\prime})$-multiplicative map $\varphi:G\rightarrow\textrm{Sym}(A)$ such that $|S^{\prime}|>(1-\epsilon^{\prime})|A|$, where we define,
@@
Fix finite subsets $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$. Let the free generators of $G$ be $\{g_{1},g_{2},\cdots\}$ (either a finite set or a sequence, depending on whether $G$ is finitely generated or not). Let $F^{-1}=\{f_{1},\cdots,f_{n}\}$. We write $f_{k}=g_{i_{k,1}}^{\epsilon_{k,1}}\cdots g_{i_{k,m_{k}}}^{\epsilon_{k,m_{k}}}$, where $\epsilon$’s are in {±$\{\pm 1\}$. Let $B\subseteq X$ be a finite set containing $E$ and all elements of $X$ of the form $\alpha(g_{i_{k,l}}^{\epsilon_{k,l}}\cdots g_{i_{k,m_{k}}}^{\epsilon_{k,m_{k}}})x$ for all $x\in E$, $1\leq k\leq n$, $1\leq l\leq m_{k}$. We define a homomorphism $\psi:G\rightarrow\textrm{Sym}(B)$ by defining, for a free generator $g_{i}$, $\psi(g_{i})$ to be any element of $\textrm{Sym}(B)$ s.t. $\psi(g_{i})b=\alpha(g_{i})b$ whenever $\alpha(g_{i})b\in B$, then extending it to a homomorphism from $G$. Let $A=\textrm{Sym}(B)$ regarded as a finite set. Then, the left multiplication action of $\textrm{Sym}(B)$ on itself gives rise to an inclusion $\textrm{Sym}(B)\hookrightarrow\textrm{Sym}(A)$. Let $\varphi:G\rightarrow\textrm{Sym}(A)$ be $\psi$ composed with this inclusion.

### v3

@@ 1. Introduction
Let $G$ be a countable discrete group, $X$ be a countable discrete set, $\alpha:G\curvearrowright X$ be an action, AA be a finite set, $\varphi:G\rightarrow\textrm{Sym}(A)$ be a map. For a finite subset $F\subseteq G$ and $\epsilon>0$, $\varphi$ is called $(F,\epsilon)$-multiplicative if $d(\varphi(gh),\varphi(g)\varphi(h))<\epsilon$ for all $g,h\in F$, where $d$ denotes the normalized Hamming distance. For finite subsets $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$, $\varphi$ is called an $(F,E,\epsilon)$-orbit approximation of $\alpha$ if there exists a finite set $B$ and a subset $S\subseteq A$ s.t. $|S|>(1-\epsilon)|A|$ and for each $s\in S$ there is an injective map $\pi_{s}:E\hookrightarrow B$ s.t. $\pi_{\varphi(g)s}(x)=\pi_{s}(\alpha(g^{-1})x)$ for all $s\in S$, $g\in F$, $x\in E$, whenever $\varphi(g)s\in S$ and $\alpha(g^{-1})x\in E$.
@@
$\alpha$ is called sofic if for all finite subsets $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$, there exists a finite set AA and a map $\varphi:G\rightarrow\textrm{Sym}(A)$ which is unital, $(F,\epsilon)$-multiplicative, and an $(F,E,\epsilon)$-orbit approximation of $\alpha$.
@@
The first example is the left action of a sofic group $G$ on itself. We show (in Theorem 2.14 and Theorem 2.12) that this is sofic if and only if $G$ is sofic. More generally, we are able to show (in Theorem 2.14) that $G$ acting on the coset space $G/H$ is sofic where $H$ is any locally finite subgroup. On the other side, we have (see Theorem 2.12) that if the action of $G$ on a coset space $G/H$ is sofic, then there is a normal subgroup NN of $G$ such that $N\leq H$ and $G/N$ is sofic. This by itself could give help to finding new examples of sofic groups.  [^1] We would like to remind the readers that it remains a puzzling question whether $G/N$ is sofic for $G$ sofic and NN is a finite normal subgroup, say even $|N|=2$.
@@
The proof of the main result is very much inspired by and pushes the ideas of [HS18]. The above result also applies to the setting of amalgamated free generalized wreath product (see Theorem 3.7). The unfamiliar reader is directed to Definition $3.$ At this point we would like to mention that the natural variant of the above result can also be proved in the context of hyperlinearity, see Theorems 3.8, 3.9 and Corollary 3.10.
@@ 2. Sofic actions
Let $G$ be a countable discrete group, $X$ be a countable discrete set, $\alpha:G\curvearrowright X$ be an action, AA be a finite set, $\varphi:G\rightarrow\textrm{Sym}(A)$ be a map (not necessarily a homomorphism):
@@
    Recall that $G$ is called sofic if for all finite subsets $F\subseteq G$ and $\epsilon>0$ there exists a finite set AA and a map $\varphi:G\to\textrm{Sym}(A)$ which is unital, $(F,\epsilon)$-multiplicative, and $d(1,\varphi(g))>1-\epsilon$ for all $g\in F\setminus\{e\}.$
@@
    $\alpha$ is called sofic if for all finite subsets $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$, there exists a finite set AA and a map $\varphi:G\rightarrow\textrm{Sym}(A)$ which is unital, $(F,\epsilon)$-multiplicative, and an $(F,E,\epsilon)$-orbit approximation of $\alpha$.
@@
Let $n]=\{1,\cdots,n\}$, $\mu_{n}$ be the normalized counting measure on $[n]$, $\mathcal{U}$ be a free ultrafilter on $\mathbb{N}$. Then the Loeb measure space, denoted by $\prod_{\mathcal{U}}([n],\mu_{n})$, is defined to have the underlying set $\prod_{\mathcal{U}}[n]$. For a sequence $(A_{n}\subseteq[n])$, we define $\mu_{\mathcal{U}}((A_{n})_{\mathcal{U}})=\lim_{\mathcal{U}}\mu_{n}(A_{n})$. This can be extended to a measure on the $\sigma$-algebra generated by all such $(A_{n})_{\mathcal{U}}$.
@@
Let $\mathcal{U}$ be a free ultrafilter on $\mathbb{N}$. For each $n$, let $X_{n}$ be the collection of all maps $n]\rightarrow\mathbb{N}$ and $d_{n}$ be the normalized Hamming distance on $X_{n}$. Let $\mathbb{X}^{2}_{\mathcal{U}}=\prod_{\mathcal{U}}(X_{n},d_{n})$. We define the following action $\mathbb{S}^{2}_{\mathcal{U}}$ of the universal sofic group $\prod_{\mathcal{U}}(S_{n},d)$ on $\mathbb{X}^{2}_{\mathcal{U}}$ by $\mathbb{S}^{2}_{\mathcal{U}}((g_{n})_{\mathcal{U}})((f_{n})_{\mathcal{U}})=(f_{n}\circ g_{n}^{-1})_{\mathcal{U}}$. We easily observe that this is a well-defined action that preserves the metric $d_{\mathcal{U}}$. We shall call $\mathbb{S}^{2}_{\mathcal{U}}$ the second universal sofic action.
@@
$(\Rightarrow)$ Fix increasing sequences of finite subsets $F_{1}\subseteq F_{2}\subseteq\cdots\subseteq G$ and $E_{1}\subseteq E_{2}\subseteq\cdots\subseteq X$ s.t. $\cup_{i}F_{i}=G$, $\cup_{i}E_{i}=X$. Fix a decreasing sequence $\epsilon_{i}>0$ s.t. $\lim_{i}\epsilon_{i}=0$. For each $i$, let $\varphi_{i}:G\rightarrow\textrm{Sym}(A_{i})$ be a unital, $(F_{i},\epsilon_{i})$-multiplicative, and an $(F_{i},E_{i},\epsilon_{i})$-orbit approximation of $\alpha$. By taking the Cartesian products of $A_{i}$ with auxiliary finite sets if necessary, we may assume $|A_{i}|$ is strictly increasing. By definition of $(F_{i},E_{i},\epsilon_{i})$-orbit approximation, there exists a finite set $B_{i}$ and a subset $S_{i}\subseteq A_{i}$ s.t. $|S_{i}|>(1-\epsilon_{i})|A_{i}|$ and for each $s\in S_{i}$ there is an injective map $\pi^{i}_{s}:E_{i}\hookrightarrow B_{i}$ s.t. $\pi^{i}_{\varphi_{i}(g)s}(x)=\pi^{i}_{s}(\alpha(g^{-1})x)$ for all $s\in S_{i}$, $g\in F_{i}$, $x\in E_{i}$, whenever $\varphi_{i}(g)s\in S_{i}$ and $\alpha(g^{-1})x\in E_{i}$. By embedding $B_{i}$ into $\mathbb{N}$ we may take the co-domain of $\pi^{i}_{s}$ to be $\mathbb{N}$. Let $\mathcal{U}$ be any free ultrafilter on $\mathbb{N}$ containing the set $\{|A_{i}|:i\geq 1\}$. We may then define $\varphi:G\rightarrow\prod_{\mathcal{U}}(S_{n},d)$ by defining $\varphi(g)=(g_{n})_{\mathcal{U}}$ with $g_{n}=\varphi_{i}(g)$ whenever $n=|A_{i}|$ and $g_{n}=1$ otherwise. Since $\{|A_{i}|:i\geq 1\}\in\mathcal{U}$ and $\varphi_{i}$ is $(F_{i},\epsilon_{i})$-multiplicative, we see that $\varphi$ is a group homomorphism.
@@
$(\Leftarrow)$ For each $g\in G$, we shall write $\varphi(g)=(g_{n})_{\mathcal{U}}$. Since $\varphi(1_{G})=1$, we shall in particular choose $\varphi(1_{G})=(1)_{\mathcal{U}}$. Let $\varphi_{n}:G\rightarrow S_{n}$ be defined by $\varphi_{n}(g)=g_{n}$. Now, fix finite $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$, we shall show that there exists $n$ s.t. $\varphi_{n}$ is unital, $(F,\epsilon)$-multiplicative, and an $(F,E,\epsilon)$-orbit approximation of $\alpha$. We observe that $\varphi_{n}$ is unital for all $n$. Since $F$ is finite and $\varphi$ is a group homomorphism, there exists $L_{1}\in\mathcal{U}$ s.t. for all $n\in L_{1}$, $\varphi_{n}$ is $(F,\epsilon)$-multiplicative.
@@
Suppose $G$ is a sofic group. Let $F\subseteq G$ be a finite subset and $\epsilon>0$. Then there exists a finite set AA and a unital, $(F,\epsilon)$-multiplicative map $\varphi:G\rightarrow\textrm{Sym}(A)$ such that $|S|>(1-\epsilon)|A|$, where we define,
@@
Since $G$ is sofic, there exists a finite set AA and a map $\varphi:G\rightarrow\textrm{Sym}(A)$ which is unital and $(F^{\prime},\epsilon^{\prime})$-multiplicative and satisfies $d(1,\varphi(g))>1-\epsilon^{\prime}$ for all non-identity $g\in F^{\prime}$.
Fix $\in F, ghg\neq h$, then, since $F$ is symmetric and since $\varphi$ is unital and $(F^{\prime},\epsilon^{\prime})$-multiplicative, we have,
@@
Hence, $|\{s\in A:\varphi(g)s\neq\varphi(h)s\}|=|\{s\in A:s\neq\varphi(g)^{-1}\varphi(h)s\}|>(1-3\epsilon^{\prime})|A|$. So,
@@
We also observe that, for any fixed $g,h\in F$, $|\{s\in A:\varphi(gh)s=\varphi(g)\varphi(h)s\}|>(1-\epsilon^{\prime})|A|$. Therefore,
@@
Let $F\subseteq G$, $E\subseteq G/N$ be finite subsets. Fix $\epsilon>0$. Let $q:G\rightarrow G/N$ be the natural quotient map. Fix $\sigma:G/N\rightarrow G$ an arbitrary section of $q$. Let $U=\{\sigma(x)^{-1}g\sigma(g^{-1}x):x\in E,g\in F\}$. We observe that $U$ is finite and $U\subseteq N$. Let $N^{\prime}=\langle U\rangle$. Since NN is locally finite, $N^{\prime}$ is a finite group. Let,
@@
We observe that since $N^{\prime}$ is a subgroup, $1_{G}\in N^{\prime}$, so $\sigma(E)^{-1}\subseteq N^{\prime}\cdot\sigma(E)^{-1}\subseteq F^{\prime}$. Let $\epsilon^{\prime}=\frac{\epsilon}{|E|+1}$. Since $F^{\prime}$ is finite, by Lemma 2.13, there is a finite set AA and a unital, $(F^{\prime},\epsilon^{\prime})$-multiplicative map $\varphi:G\rightarrow\textrm{Sym}(A)$ such that $|S^{\prime}|>(1-\epsilon^{\prime})|A|$, where we define,
@@
Fix finite subsets $F\subseteq G$, $E\subseteq X$, and $\epsilon>0$. Let the free generators of $G$ be $\{g_{1},g_{2},\cdots\}$ (either a finite set or a sequence, depending on whether $G$ is finitely generated or not). Let $F^{-1}=\{f_{1},\cdots,f_{n}\}$. We write $f_{k}=g_{i_{k,1}}^{\epsilon_{k,1}}\cdots g_{i_{k,m_{k}}}^{\epsilon_{k,m_{k}}}$, where $\epsilon$’s are in $\{\pm 1\}$. Let $B\subseteq X$ be a finite set containing $E$ and all elements of $X$ of the form $\alpha(g_{i_{k,l}}^{\epsilon_{k,l}}\cdots g_{i_{k,m_{k}}}^{\epsilon_{k,m_{k}}})x$ for all $x\in E$, $1\leq k\leq n$, $1\leq l\leq m_{k}$. We define a homomorphism $\psi:G\rightarrow\textrm{Sym}(B)$ by defining, for a free generator $g_{i}$, $\psi(g_{i})$ to be any element of $\textrm{Sym}(B)$ s.t. $\psi(g_{i})b=\alpha(g_{i})b$ whenever $\alpha(g_{i})b\in B$, then extending it to a homomorphism from $G$. Let $A=\textrm{Sym}(B)$ regarded as a finite set. Then, the left multiplication action of $\textrm{Sym}(B)$ on itself gives rise to an inclusion $\textrm{Sym}(B)\hookrightarrow\textrm{Sym}(A)$. Let $\varphi:G\rightarrow\textrm{Sym}(A)$ be $\psi$ composed with this inclusion.